jagr2713 (jagr2713):

-x-3y=-11 y=8x-13 Solve by substitution

jagr2713 (jagr2713):

@nincompoop @Nnesha

OpenStudy (anonymous):

wow this is tricky but i think i can figure it out

jagr2713 (jagr2713):

Yes I tricking me

Nnesha (nnesha):

well 2nd equation is already solved for y so substitute y for 2nd equation into -x-3y=-11

jagr2713 (jagr2713):

ohhhh so we just solve for the fist equation? so we get 3y=x-11

Nnesha (nnesha):

for substitution you need to solve one of the equation for x `or` y but 2nd equation is already solved for y then just plug that into the first equation

jagr2713 (jagr2713):

so just plug the 2nd and put for x above or?

Nnesha (nnesha):

y=8x -13 so you can replace y with 8x-13 \[\huge\rm -x-3\color{ReD}{y}=-11\]

Nnesha (nnesha):

y=8x-13 so replace y not x

jagr2713 (jagr2713):

ohoho so we get -x-3(8x-13)=-11 -x-24x+39=-11 -25x+39=-11 -25x=-50 x=2

jagr2713 (jagr2713):

then now we plug it into y=8x-13 y=8(2)-13 y=3

Nnesha (nnesha):

perfect!

jagr2713 (jagr2713):

wow that was easy I just got lost because it came with an extra number :D

Nnesha (nnesha):

wait a sec hmm

jagr2713 (jagr2713):

I got it correct no need :D

Nnesha (nnesha):

lol i was just checking

Nnesha (nnesha):

good job! :D